A valószínűségszámítás I. gyakorlat április 2. dolgozatának eredményei.

dolgozat példák:  1.  2.  3.  4.  5.  6.

diákok:

Barabás Timea  0  0  0  0  4  0
Csaba Anikó  4  3  4  2  4  1
Dorogi Edina  4  1  0  0  1  5
Jámbor Tamás  4  0  1  1  4  4
Hubert Balázs  0  0  0  0  4  2
Medve Diana  0  0  4  0  4  4
Mészáros Fruzsina  4  2  0  4  4  5
Öldi Kinga  4  1  0  0  0  1
Suta Judit  4  1  1  0  4  3
Tanka Irén  0  1  0  0  4  1